You decide to take advantage of the current online dating craze and start your own Web site. You know that you have 450 people who will sign up immediately and, through a careful marketing research and analysis, determine that membership can grow by 27 percent in the first two years, 22 percent in year 3, and 18 percent in Year 4. How many members do you expect to have at the end of four years? (Round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest whole number.)
Answer #1

Number of members in every year need to be calculated.

Starting member numbers = 450

In Year 1, members = 450 * (1 + 27%) = 571.500000

In Year 2, members = 572 * (1 + 27%) = 725.805000

In Year 3, members = 726 * (1 + 22%) = 885.482100

In Year 4, members = 885 * (1 + 18%) = 1044.868878

Answer = 1,045